# Doubly-Robust LLM-as-a-Judge: Externally Valid Estimation with Imperfect Personas

### v1

@@ A.4 Statistical Frameworks for Sample Efficient-Estimation
Recent developments in black-box predictive models that can operate on multi-modal representations has spurred significant interest in how these predictions might be used to improve statistical inference (Angelopoulos et al., 2023a; b; Fisch et al., 2024; Eyre & Madras, 2024; Dorner et al., 2024; Ji et al., 2025; Saad-Falcon et al., 2023; Fogliato et al., 2024). These frameworks address the challenge of making valid statistical inferences when labeled data is scarce but black-box predictions cheap and abundant. We briefly review developments in this literature before identifying key gaps addressed by our approach.
@@
Prediction-Powered Inference (PPI) uses predictions from a black-box machine learning model to tighten confidence intervals when labeled data is scarce (Angelopoulos et al., 2023a; b). This is done through the addition of a “rectifier” — a mean zero term that contrasts the performance of the model’s predictions on the labeled and unlabeled points. While initial variants of PPI were not computationally efficient, Angelopoulos et al. (2023b) introduce a PPI++ framework, which introduces a “trust” parameter $\lambda$ to control the magnitude of the rectifier. This allows for efficiently computable confidence sets that are provably tighter than those computed just from labeled data. We also emphasize recent work due to Ji et al. (2025), which shows that PPI++ is just a special parametric class of solutions for M-estimation with surrogates outcomes — a classical, well-studied problem in the causal inference/missing data literature (Robins et al., 1994; Robins & Rotnitzky, 1995; Tsiatis, 2006). While the authors do not directly mention Neyman orthogonal scores (Neyman, 1979; Chernozhukov et al., 2018) in their work, they construct what is implicitly a Neyman orthogonal score for the problem at hand and propose a solution based on cross-fitting. Additional theoretical developments along these lines have been proposed — Ao et al. (2024) propose a framework for adaptive estimation of linear functionals based on supplied predictions. Likewise, Xu et al. (2025) consider a general semi-parametric framework for estimating functionals of the data generating distribution in the presence of ML predictions. However, none of the aforementioned works provides a framework usable in settings where (a) unlabeled and labeled samples come from different distributions (i.e. covariate shift) and (b) data is missing at random (MAR), (i.e. the probability that outcomes are observed for any given individual depend on their features). We close this gap by proposing a doubly-robust estimator (Bang & Robins, 2005) and a general algorithm based on cross-fitting (Chernozhukov et al., 2018) for solving M-estimation problems in the presence of both covariate shift and heterogeneity in data missingness. We also incorporate recent ideas on “Riesz losses” (Chernozhukov et al., 2022b; 2023; a; Hirshberg & Wager, 2021), loss functions that specify complicated nuisance functions as their minimizers. By using Riesz losses to learn the re-weighting function $\alpha_{0}(W,C)$, we avoid constructing plug-in estimates for $\omega_{0}$ and $\pi_{0}$ and computing their quotient, which can result in high bias.
@@ C.1 Cross-Fitting for M-Estimators
To show $R_{1}=o_{\mathbb{P}}(1)$, it suffices to show that $\sup_{\theta\in\Theta}\|J_{N_{s}}(\theta,\widehat{\alpha})-J(\theta,\widehat{\alpha})\|_{op}=o_{\mathbb{P}}(1)$. We know that for any fixed square-integrable function $\alpha(w,c)$, since $\nabla_{\theta}m(w,y;\theta)$ is bounded above in operator norm by some constant $D$, we have $\|\alpha(w,c)\nabla_{\theta}m(w,y;\theta)\|_{op}\leq D|\alpha(w,c)|$, and so the collection of scores possesses an integrable envelope. Further, since $\nabla_{\theta}m(w,y;\theta)$ is continuous in $\theta$, the score $\alpha(w,c)\nabla_{\theta}m(w,y;\theta)$ is continuous as well. Lastly, since $\Theta$ is compact, Lemma 2.4 of Newey & McFadden (1994) yields that $\{\alpha(w,c)\nabla_{\theta}m(w,y;\theta):\theta\in\Theta\}$ is a weak Glivenko-Cantelli class, i.e. that
@@
both converge to zero in probability. Since we have assumed $m(w,y;\theta)$ is bounded by assumption, we can again use Lemma 2.4 of Newey & McFadden (1994) to obtain that $\sup_{\theta}\|M_{N_{s}}(\theta,\alpha)-M(\theta,\alpha)\|=o_{\mathbb{P}}(1)$ for each fixed, square-integrable $\alpha$. The bounded convergence theorem them yields that, for any $\epsilon>0$,
@@ Appendix D Details on Riesz Losses
In this section, we discuss the Riesz loss outlined in Equation (4). Introduced in Chernozhukov et al. (2022b) and later in expanded upon in Chernozhukov et al. (2022a; 2023), Riesz losses provide a principled approach rooted in empirical risk minimization framework for estimating complicated nuisances.
@@ E.1 Synthetic Dataset
$\hat{\mu}(X)$ $:=\rho\cdot Y+\sqrt{1-\rho^{2}}\cdot\epsilon+b,\text{ for }\epsilon\sim N(0,\sigma^{2}_{Y}),b\in\mathbb{R}$ (14)
where $\Phi(X)$ represents polynomial feature expansion with interactions (degree 2), $p_{s,j}$ and $p_{t,j}$ are the Bernoulli parameters for feature $j$ in source and target domains respectively, $\rho$ controls the correlation between true and surrogate ratings, and $b$ represents surrogate bias.
@@
    $\hat{Y}=\text{clip}(\rho\cdot Y+\sqrt{1-\rho^{2}}\cdot Z+b,y_{\min},y_{\max})$
@@ E.4 Persona Simulation Framework
To simulate covariate shifts that may occur in real-world settings, we reference statistics reported by the U.S. Census Bureau (Guzman & Kollar, 2023) and the rater demographic distribution already present in DICES (Aroyo et al., 2023) which are reported in table 2.
Table 2: Population statistics used to define source and target rater distributions. Source distributions $P_{s}(X)$ are based on DICES-reported rater characteristics, while target distributions $P_{t}(X)$ follow U.S. Census Bureau statistics (Guzman & Kollar, 2023).
@@ E.4.1 LLM Usage
We use two LLMs (GPT-5 and GPT4o-Mini) to simulate human ratings for PRISM and four models to produce persona-based surrogate ratings for DICES (Claude-Sonnet-3.5, Claude-Haiku-3.5, GPT4o-Mini, and GPT-5). We use API calls for all LLMs with temperatures of 0.7 (for results reported in the main paper) and 0.9 for Anthropic models and 1 for OpenAI models. We use GPT-5 with reasoning set to “low.” Our exact system and user prompts are shown in Fig6,7,8.

### v2

@@ A.4 Statistical Frameworks for Sample Efficient-Estimation
Recent developments in black-box predictive models that can operate on multi-modal representations has spurred significant interest in how these predictions might be used to improve statistical inference (Byun et al., 2025; Angelopoulos et al., 2023a; b; Fisch et al., 2024; Eyre and Madras, 2024; Dorner et al., 2024; Ji et al., 2025; Saad-Falcon et al., 2023; Fogliato et al., 2024). These frameworks address the challenge of making valid statistical inferences when labeled data is scarce but black-box predictions cheap and abundant. We briefly review developments in this literature before identifying key gaps addressed by our approach.
@@
Prediction-Powered Inference (PPI) uses predictions from a black-box machine learning model to tighten confidence intervals when labeled data is scarce (Angelopoulos et al., 2023a; b). This is done through the addition of a “rectifier” — a mean zero term that contrasts the performance of the model’s predictions on the labeled and unlabeled points. While initial variants of PPI were not computationally efficient, Angelopoulos et al. (2023b) introduce a PPI++ framework, which introduces a “trust” parameter $\lambda$ to control the magnitude of the rectifier. This allows for efficiently computable confidence sets that are provably tighter than those computed just from labeled data. We also emphasize recent work due to Ji et al. (2025), which shows that PPI++ is just a special parametric class of solutions for M-estimation with surrogates outcomes — a classical, well-studied problem in the causal inference/missing data literature (Robins et al., 1994; Robins and Rotnitzky, 1995; Tsiatis, 2006). While the authors do not directly mention Neyman orthogonal scores (Neyman, 1979; Chernozhukov et al., 2018) in their work, they construct what is implicitly a Neyman orthogonal score for the problem at hand and propose a solution based on cross-fitting. Additional theoretical developments along these lines have been proposed — Ao et al. (2024) propose a framework for adaptive estimation of linear functionals based on supplied predictions. Likewise, Xu et al. (2025) consider a general semi-parametric framework for estimating functionals of the data generating distribution in the presence of ML predictions.
However, none of the aforementioned works provides a framework usable in settings where (a) unlabeled and labeled samples come from different distributions (i.e. covariate shift) and (b) data is missing at random (MAR), (i.e. the probability that outcomes are observed for any given individual depend on their features). [^7] Concurrent work by Ulichney and Coston (2025) also studies model evaluation under joint covariate shift and selection bias, deriving a bias-corrected estimator based on influence functions. While their focus is on pre-deployment risk estimation for black-box prediction models, our framework targets M-estimation in the context of GenAI evaluation. We additionally leverages persona ratings as surrogate outcomes to improve inference. We close this gap by proposing a doubly-robust estimator (Bang and Robins, 2005) and a general algorithm based on cross-fitting (Chernozhukov et al., 2018) for solving M-estimation problems in the presence of both covariate shift and heterogeneity in data missingness. We also incorporate recent ideas on “Riesz losses” (Chernozhukov et al., 2022a; 2023; b; Hirshberg and Wager, 2021), loss functions that specify complicated nuisance functions as their minimizers. By using Riesz losses to learn the re-weighting function $\alpha_{0}(W,C)$, we avoid constructing plug-in estimates for $\omega_{0}$ and $\pi_{0}$ and computing their quotient, which can result in high bias.
@@ C.1 Cross-Fitting for M-Estimators
To show $R_{1}=o_{\mathbb{P}}(1)$, it suffices to show that $\sup_{\theta\in\Theta}\|J_{N_{s}}(\theta,\widehat{\alpha})-J(\theta,\widehat{\alpha})\|_{op}=o_{\mathbb{P}}(1)$. We know that for any fixed square-integrable function $\alpha(w,c)$, since $\nabla_{\theta}m(w,y;\theta)$ is bounded above in operator norm by some constant $D$, we have $\|\alpha(w,c)\nabla_{\theta}m(w,y;\theta)\|_{op}\leq D|\alpha(w,c)|$, and so the collection of scores possesses an integrable envelope. Further, since $\nabla_{\theta}m(w,y;\theta)$ is continuous in $\theta$, the score $\alpha(w,c)\nabla_{\theta}m(w,y;\theta)$ is continuous as well. Lastly, since $\Theta$ is compact, Lemma 2.4 of Newey and McFadden (1994) yields that $\{\alpha(w,c)\nabla_{\theta}m(w,y;\theta):\theta\in\Theta\}$ is a weak Glivenko-Cantelli class, i.e. that
@@
both converge to zero in probability. Since we have assumed $m(w,y;\theta)$ is bounded by assumption, we can again use Lemma 2.4 of Newey and McFadden (1994) to obtain that $\sup_{\theta}\|M_{N_{s}}(\theta,\alpha)-M(\theta,\alpha)\|=o_{\mathbb{P}}(1)$ for each fixed, square-integrable $\alpha$. The bounded convergence theorem them yields that, for any $\epsilon>0$,
@@ Appendix D Details on Riesz Losses
In this section, we discuss the Riesz loss outlined in Equation (4). Introduced in Chernozhukov et al. (2022a) and later in expanded upon in Chernozhukov et al. (2022b; 2023), Riesz losses provide a principled approach rooted in empirical risk minimization framework for estimating complicated nuisances.
@@ E.1 Synthetic Dataset
$\hat{\mu}(X)$ $:=\text{clip}(\rho\cdot Y+\sqrt{1-\rho^{2}}\cdot Z\sigma_{Y}+\eta(y_{\max}-y_{\min}),y_{\min},y_{\max})$ (13)
where $\Phi(X)$ represents polynomial feature expansion with interactions (degree 2), $p_{s,j}$ and $p_{t,j}$ are the Bernoulli parameters for feature $j$ in source and target domains respectively, $\rho$ controls the correlation between true and surrogate ratings, $y_{\min},y_{\max}$ define the interval of the rating scale, and $Z$ represents independent Gaussian noise.
@@
    $\hat{Y}=\text{clip}(\rho\cdot Y+\sqrt{1-\rho^{2}}\cdot Z\sigma_{Y}+\eta(y_{\max}-y_{\min}),y_{\min},y_{\max})$
@@ E.4 Persona Simulation Framework
To simulate covariate shifts that may occur in real-world settings, we reference statistics reported by the U.S. Census Bureau (Guzman and Kollar, 2023) and the rater demographic distribution already present in DICES (Aroyo et al., 2023) which are reported in table 3.
Table 3: Population statistics used to define source and target rater distributions. Source distributions $P_{s}(X)$ are based on DICES-reported rater characteristics, while target distributions $P_{t}(X)$ follow U.S. Census Bureau statistics (Guzman and Kollar, 2023).
@@
Figure 6: Visualizing Sources of Covariate Shift in the DICES Dataset. (a) Probability of sampling two demographic subgroups as a function of rater shift magnitude. At low rater shift (magnitude = 0), Asian Gen Z college graduates are more likely to be sampled; at high rater shift (magnitude $\geq 0.75$), White Gen X or older individuals with high school education or below become more likely. (b) Sinkhorn Distance between source and target distributions as a function of source/target content overlap $\zeta$ (x-axis) and rater shift magnitude (y-axis). The vertical orientation of contour lines indicates that content features have a larger impact on Sinkhorn Distance than the rater features. (c) Examples showing how sampling probabilities of source and target samples vary with $\zeta$. The top comment (Religious + Misinformation harms) is from the target distribution and becomes more likely to be included in source data as $\zeta\rightarrow 1$ (purple cone). Conversely, the bottom comment (Regulated Goods) is from the source distribution and becomes less likely to be retained as $\zeta\rightarrow 1$ (green cone). Together, these panels illustrate that content covariate shift plays a larger role than rater covariate shift in the DICES dataset.
Figure 7: Average correlation between each embedding feature and human ratings (Y) and each embedding feature and source/target membership (A). Average is over all features recovered in the projected sub-space of a specified embedding dimension. We observe an “elbow” shape in the plot, where the correlation diminishes after each additional feature beyond 15.
@@ E.4.1 LLM Usage
We use two LLMs (GPT-5 and GPT4o-Mini) to simulate human ratings for PRISM and four models to produce persona-based surrogate ratings for DICES (Claude-Sonnet-3.5, Claude-Haiku-3.5, GPT4o-Mini, and GPT-5). We use API calls for all LLMs with temperatures of 0.7 (for results reported in the main paper) and 0.9 for Anthropic models and 1 for OpenAI models. We use GPT-5 with reasoning set to “low.” Our exact system and user prompts are shown in Fig8,9,10.